# SaraCoder: Orchestrating Semantic and Structural Cues for Profit-Oriented Repository-Level Code Completion

### v1

@@ Introduction
Prompt Generation
@@
Experiments
Experimental Settings
@@ 3. RQ3
    Synergistic Gain Property: How does SaraCoder perform when integrated orthogonally with other cross-file methods?
@@ 4. RQ4
Datasets
We primarily utilize two datasets here: CrossCodeEval and RepoEval-Updated. Table 1 shows details.
  • CrossCodeEval (Ding et al. 2023): This benchmark evaluates code completion in complex Cross-File scenarios like type inference and dependency analysis. It is ideal for assessing performance that requires a deep understanding of code across multiple files.
  • RepoEval-Updated (Liu et al. 2024): Expanded from RepoEval (2023), this new version, includes repositories of varying scales and duplication rates, offering a better way to evaluate In-File completion performance.
We use CrossCodeEval to test models on code completion tasks that involve complex cross-file dependencies. RepoEval-Updated will assess other aspects like basic syntax, common API usage, and local context understanding.
CrossCodeEval RepoEval-Updated
Python Java Python Java
Total Repositories 471 239 10 8
Total Files 1368 745 3258 8260
Total Task cases 2665 2139 2000 1600
Applicable Scenarios Cross-File completion In-File completion
Table 1: CrossCodeEval vs. RepoEval-Updated comparison.
Evaluation Indicators Setting
In this study, the following several evaluation indicators are used to assess the effect of code completion (Lu et al. 2021; Ding et al. 2023).
  • Identifier Exact Match (ID_EM): The percentage of identifiers (variables, functions, etc.) perfectly matching the ground code. A high ID_EM means the model correctly understands the context and chooses the right names.
  • Identifier F1 Score (ID_F1): A more nuanced evaluation of identifier matching by combining precision and recall. This integration offers a more comprehensive assessment of identifier completion quality.
Baseline Setting
We employ the following five methods as controls to evaluate the effectiveness of retrieval-augmented generation (RAG) in code completion: No RAG (Zero-shot-only baseline), Shifted RAG (Target-context dynamic retrieval), Vanilla RAG (Exemplar-similarity fixed retrieval), Repocoder (Iterative-fragment integration (Zhang et al. 2023)), Graphcoder (Structure-modeling CCG utilization (Liu et al. 2024)).
@@
Model Selection
In this experiment, we select Codegen2-7b, Codegen25-7b (Nijkamp et al. 2023b, a), CodeLlama-7b-Instruct (Rozière et al. 2024) and deepseek-coder-6.7b-instruct (Guo et al. 2024) for code completion task inference.
@@
Table 3: Performance benefits of SaraCoder when integrated orthogonally with other cross-file approaches ($\%$).
@@
Main Results
For RQ1: Dominant Cross-File Code Accuracy.
Table 2 illustrates that SaraCoder surpasses the top-performing Repocoder on the CrossCodeEval dataset, achieving an average improvement of $1.50$ in EM, $0.77$ in ES, $1.11$ in ID_EM, and $0.61$ in ID_F1. This indicates SaraCoder provides more effective information and generates code with higher semantic accuracy, better capturing intended functionality. The enhanced ID_EM further shows SaraCoder’s superior ability to interpret context and select appropriate identifiers. These advancements effectively mitigate misleading superficial similarity and external symbol ambiguity, leading to more reliable and contextually relevant code. For Java code completion, SaraCoder shows better EM and ID_EM, with slightly lower ES and F1 scores. Because of Java’s static typing and verbosity, even small structural changes (e.g., bracket placement) can heavily affect ES/F1 scores, despite functional correctness.
For RQ2: Cost-Optimized Accuracy Advantage In Cross-File.
We experiment with code completion efficiency using codegen25-7b and Graphcoder. Our goal is to see how retrieving more similar cases (top_ktop\_k) impacts accuracy. Since using fewer top_ktop\_k cases saves input tokens [^2] For relevant explanations, please refer to the Appendix D.2, this study shows the balance between resources and accuracy. Our experiments demonstrate significant performance saturation for both retrieval methods when top_ktop\_k reaches 3-4, with no observable fluctuations upon increasing to top_k=10top\_k=10. SaraCoder achieves comprehensive superiority in Python tasks (e.g., 9.4% EM improvement) while maintaining advantages in Java tasks despite a marginal $0.1$ decrease in ES. Crucially, under resource-constrained top_k=1top\_k=1 conditions: all Python metrics outperform the baseline; three Java metrics (EM/ES/ID_EM) show improvements; and Java ID_F1 initially trails ($35.22$ vs. $35.27$) but ultimately surpasses the baseline at saturation ($35.84$ vs. $35.77$). Our method achieves performance breakthroughs at lower computational cost (stable at $top\_k\approx 4$) by reducing redundant and homogeneous cases (Figure 4).
For RQ3: Synergistic Integration of SaraCoder Achieves Enhanced Completion.
We examine two prominent methods that demonstrate exceptional performance in Cross-File scenarios. (1) Repocoder (Zhang et al. 2023), distinct from the original, assumes that if code snippets are similar, their subsequent content is also likely relevant. In the next search round, it specifically gets the code following those similar snippets (hereafter referred to as Repocoder). (2) Draco (Cheng, Wu, and Hu 2024), analyzes code to create entity dependency graphs, allowing detailed background knowledge retrieval. It then uses this information to create structured prompts. Currently, Draco only works with Python. As shown in Table 3, adding our method significantly boosts all four Python metrics (by $3.42$ to $4.52$) compared to using Repocoder or Draco alone. For Java, our method improves EM by 0.45 and ID_EM by $0.33$ over Repocoder, showing SaraCoder exhibits significant synergistic gain property with existing cross-file methods.  [^3] You can find the causes of synergistic gains in Appendix D.3.
Figure 5: Ablation study. (Each three-data-point group represents CodeGen2-7B, CodeGen2.5-7B, and CodeLlama-7B-Instruct models. Bar lengths show their average performance, with I-shaped error bars indicating S.D.)
For RQ4: Enhanced In-File Accuracy and Resource Efficiency.
On the RepoEval-Updated dataset (Table 4), SaraCoder shows superior semantic and identifier accuracy (surpassing the top-performing Graphcoder: +0.547+0.547 EM, +0.737+0.737 ES, +0.125+0.125 ID_EM, and +0.667+0.667 F1) for both Python and Java code completion. The cost analysis (Appendix D.1) further indicates SaraCoder generally performs better and exhibits higher stability across most Python metrics (excluding EM) and all Java metrics. This makes it particularly effective for resource-constrained environments, especially at lower top_ktop\_k values. However, SaraCoder’s gains over Graphcoder in code and identifier matching are smaller here than on CrossCodeEval. This is primarily because RepoEval-Updated projects contain a higher prevalence of similar code snippets, resulting in reduced code diversity. Overall, the conclusions align with those from the CrossCodeEval dataset.
Ablation Study

### v2

@@ 3.3 External-Aware Identifier Disambiguator (EAID)
Figure 3: Prompt template.
3.4 Prompt Generation
@@
4 Experiments
4.1 Experimental Settings
4.1.1 Datasets
Table 1: CrossCodeEval vs. RepoEval-Updated comparison.
CrossCodeEval RepoEval-Updated
Python Java Python Java
Total Repositories 471 239 10 8
Total Files 14348 5868 3258 8260
Total Task cases 2665 2139 2000 1600
Applicable Scenarios Cross-File completion In-File completion
We primarily utilize two datasets here: CrossCodeEval and RepoEval-Updated. (1) CrossCodeEval Ding et al. (2023): This benchmark evaluates code completion in complex Cross-File scenarios like type inference and dependency analysis. It is ideal for assessing performance that requires a deep understanding of code across multiple files. (2) RepoEval-Updated Liu et al. (2024b): Expanded from RepoEval Zhang et al. (2023), this new version, includes repositories of varying scales, offering a better way to evaluate In-File completion performance. We use CrossCodeEval to test models on code completion tasks required complex cross-file dependencies. RepoEval-Updated assesses basic syntax, common API usage, and local context understanding. Table 1 shows details.
4.1.2 Evaluation Indicators Setting
In this study, the following several evaluation indicators are used to assess the effect of code completion Lu et al. (2021); Ding et al. (2023).
  • Identifier Exact Match (ID_EM): The percentage of identifiers (variables, functions, etc.) perfectly matching the ground code. A high ID_EM score indicates the model’s strong contextual understanding, enabling it to accurately predict and generate contextually appropriate identifiers.
  • Identifier F1 Score (ID_F1): A more nuanced evaluation of identifier matching by combining precision and recall. It offers a more comprehensive assessment of identifier completion quality, particularly beneficial in scenarios where models might generate partial but correct identifier sets.
4.1.3 Baseline Setting
We employ the following five methods as controls to evaluate the effectiveness of retrieval-augmented generation (RAG) in code completion: No RAG (Zero-shot-only baseline), Shifted RAG (Target-context dynamic retrieval), Vanilla RAG (Exemplar-similarity fixed retrieval), Repocoder (Iterative-fragment integration Zhang et al. (2023)), Graphcoder (Structure-modeling CCG utilization Liu et al. (2024b)).
4.1.4 Model Selection
In this experiment, we select Codegen2-7b, Codegen25-7b Nijkamp et al. (2023b; a), CodeLlama-7b-Instruct Rozière et al. (2024) and deepseek-coder-6.7b-instruct Guo et al. (2024) for code completion task inference.
4.2 Main Results
@@ 3. RQ3
    Synergistic Gain Property: How does SaraCoder perform when integrated orthogonally with other methods that provide cross-file context?
@@ 4. RQ4
4.2.1 For RQ1: Dominant Cross-File Code Accuracy.
Table 2 illustrates that SaraCoder surpasses the top-performing Repocoder on the CrossCodeEval dataset, achieving an average improvement of $1.50$ in EM, $0.77$ in ES, $1.11$ in ID_EM, and $0.61$ in ID_F1. This indicates SaraCoder provides more effective information and generates code with higher semantic accuracy, better capturing intended functionality. The enhanced ID_EM further shows SaraCoder’s superior ability to interpret context and select appropriate identifiers. These advancements effectively mitigate misleading superficial similarity and external symbol ambiguity, leading to more reliable and contextually relevant code. For Java code completion, SaraCoder shows better EM and ID_EM, with slightly lower ES and F1 scores. This discrepancy is attributed to the inherent characteristics of Java’s static typing system and complex code structure. These features lead to the generation of code that is logically correct but contains numerous textual variations and boilerplate redundancies. Consequently, small structural deviations (e.g., misplaced brackets) are more readily penalized by metrics such as ES and F1.
4.2.2 For RQ2: Cost-Optimized Accuracy Advantage In Cross-File.
We experiment with code completion efficiency using codegen25-7b and Graphcoder. Our goal is to see how retrieving more similar cases (top_ktop\_k) impacts accuracy. Since using fewer top_ktop\_k cases saves input tokens [^2] For relevant explanations, please refer to the Appendix A.6.2, this study shows the balance between resources and accuracy. Our experiments demonstrate significant performance saturation for both retrieval methods when top_ktop\_k reaches 3-4, with no observable fluctuations upon increasing to top_k=10top\_k=10. SaraCoder achieves comprehensive superiority in Python tasks (e.g., 9.4% EM improvement) while maintaining advantages in Java tasks despite a marginal $0.1$ decrease in ES. Crucially, under resource-constrained top_k=1top\_k=1 conditions: all Python metrics outperform the baseline; three Java metrics (EM/ES/ID_EM) show improvements; and Java ID_F1 initially trails ($35.22$ vs. $35.27$) but ultimately surpasses the baseline at saturation ($35.84$ vs. $35.77$). Our method achieves performance breakthroughs at lower computational cost (stable at $top\_k\approx 4$) by reducing redundant and homogeneous cases (Figure 4).
@@
4.2.3 For RQ3: Synergistic Integration of SaraCoder Achieves Enhanced Completion.
We examine two prominent methods that demonstrate exceptional performance in Cross-File scenarios. (1) Repocoder Zhang et al. (2023), distinct from the original, assumes that if code snippets are similar, their subsequent content is also likely relevant. In the next search round, it specifically gets the code following those similar snippets (hereafter referred to as Repocoder). (2) Draco Cheng et al. (2024), analyzes code to create entity dependency graphs, allowing detailed background knowledge retrieval. It then uses this information to create structured prompts. Currently, Draco only works with Python. As shown in Table 3, adding our method significantly boosts all four Python metrics (by $3.42$ to $4.52$) compared to using Repocoder or Draco alone. For Java, our method improves EM by 0.45 and ID_EM by $0.33$ over Repocoder, showing SaraCoder exhibits significant synergistic gain property with existing cross-file methods.  [^3] You can find the causes of synergistic gains in Appendix A.6.3.
Table 3: Performance benefits of SaraCoder when integrated orthogonally with other cross-file approaches ($\%$). The CrossCodeEval dataset is used in this part.
@@
4.2.4 For RQ4: Enhanced In-File Accuracy and Resource Efficiency.
On the RepoEval-Updated dataset (Table 4), SaraCoder shows superior semantic and identifier accuracy (surpassing the top-performing Graphcoder: +0.547+0.547 EM, +0.737+0.737 ES, +0.125+0.125 ID_EM, and +0.667+0.667 F1) for both Python and Java code completion. The cost analysis (Appendix A.6.1) further indicates SaraCoder generally performs better and exhibits higher stability across most Python metrics (excluding EM) and all Java metrics. This makes it particularly effective for resource-constrained environments, especially at lower top_ktop\_k values. However, SaraCoder’s gains over Graphcoder in code and identifier matching are smaller here than on CrossCodeEval. This is primarily because RepoEval-Updated projects contain a higher prevalence of similar code snippets, resulting in reduced code diversity within the repository. Overall, the conclusions align with those from the CrossCodeEval dataset.
@@
4.3 Ablation Study